# secReferences

### v1

@@ Abstract
The kidney paired donation (KPD) program provides an innovative solution to overcome incompatibility challenges in kidney transplants by matching incompatible donor-patient pairs and facilitating kidney exchanges. To address unequal access to transplant opportunities, there are two widely used fairness criteria: group fairness and individual fairness. However, these criteria do not consider protected patient features, which refer to characteristics legally or ethically recognized as needing protection from discrimination, such as race and gender. Motivated by the calibration principle in machine learning, we introduce a new fairness criterion: the matching outcome should be conditionally independent of the protected feature, given the sensitization level. We integrate this fairness criterion as a constraint within the KPD optimization framework and propose a computationally efficient solution. Theoretically, we analyze the associated price of fairness using random graph models. Empirically, we compare our fairness criterion with group fairness and individual fairness through both simulations and a real-data example.
Keywords: calibration, integer programming, kidney paired donation, price of fairness, random graph.
@@ 1.4 Our contributions
This paper makes several contributions to the field of kidney exchange and fairness in allocation. First, we propose a new fairness criterion based on a protected feature, which has not been explored in the kidney exchange literature. We integrate this fairness criterion as a constraint within the optimization framework commonly used in kidney paired donation (KPD) programs. This flexible structure can accommodate other fairness criteria and potential recourse strategies. Furthermore, we propose a computationally efficient solution to the resulting optimization problem.
@@ 2.1.1 Classical formulation
We represent a KPD program as a directed graph $G=(V,E)$, where $V=\{v_{1},...,v_{N}\}$ denotes the vertex set and $E$ denotes the edge set. The vertex set $V$ is the set of $N$ incompatible donor-patient pairs. The edge set $E$ consists of all ordered pairs $(v_{i},v_{j})$ that the donor in vertex $v_{i}\in V$ is compatible with the patient in vertex $v_{j}\in V_{p}$. An exchange cycle, or simply cycle, is defined as a sequence of distinct incompatible donor-patient pairs. We denote a cycle $c$ as an ordered sequence of vertices $[c_{1},...,c_{|c|}]$ in $V$, where $|c|$ is the cycle length of $c$, satisfying that $(c_{i},c_{i+1})\in E$ for $11\leq i\leq|c|-$ 1, and $(c_{|c|},c_{1})\in E$. To execute the cycle, the patient in $c_{i+1}$ will receive the donor kidney of $c_{i}$ for $11\leq i\leq k-$ 1, and the patient $c_{1}$ will receive the donor kidney of $c_{k}$. Figure 1 illustrates how a two-way cycle and a three-way cycle work in KPD programs. An exchange plan is a collection of vertex-disjoint cycles in the graph.
@@
In fielded KPD programs, we usually restrict cycles to be no longer than 3. We use $\mathcal{C}$ to denote the set of all cycles under the restrictions on the length of the cycles. We assign a utility $u_{v_{i},v_{j}}$ to each edge $(v_{i},v_{j})\in E$. Based on the edge utility, we define the cycle utility $u(c)=\sum_{i=1}^{|c|-1}u_{c_{i},c_{i+1}}+u_{c_{|c|},c_{1}}$. In a classical KPD program, we aim to select disjoint packing of cycles in $\mathcal{C}$ with the maximum sum of utilities, which can be formulated as the following integer programming:
@@ 2.2.2 Individual fairness
The individual fairness utilizes a randomization policy to ensure that every patient has a similar chance to be matched. Let $\mathcal{F}\subseteq 2^{\mathcal{C}}$ be the set of possible exchange plans that consist of disjoint cycles in $\mathcal{C}$. We assign a probability distribution $\delta$ on $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Let $x_{c}(\delta)$ be the probability that the cycle $c\in\mathcal{C}$ is selected as one of the disjoints subsets. Based on $x_{c}(\delta)$, we can compute the probability of each vertex $v$ to be selected $\delta_{v}=\sum_{c\in\mathcal{C}}x_{c}(\delta)1_{v\in c}$, which is a central quantity in defining individual fairness and our new fairness in Section 3. Recall that $x_{c}$ in (1) and (2) is a binary variable indicating whether $c$ is selected or not, but here $x_{c}(\delta)$ is a continuous variable bounded between 0 and 1. Following Farnadi et al. (2021), We formulate the individual fairness as a constraint in the optimization problem below
$\max_{\delta}\quad\sum_{c\in\mathcal{C}}x_{c}(\delta)u(c)\quad\text{subject to}\quad\sum_{v\in V}|\delta_{v}-\overline{\delta}|^{p}\geq\beta^{p}.$ (3)
Specifically, the individual fairness promotes a similar chance of being selected for each patient. We choose the $L_{p}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ to measure the variation of the selection probability among all the patients, where $\overline{\delta}$ is the average selection probability of all patients. We consider a fairness constraint that the $L_{p}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ is no greater than some threshold $\beta$.
@@
3.1 Fairness formulation and algorithm
@@
In this section, we introduce a new fairness criterion based on a protected feature $A$ of patients in $V$, motivated by the concept of calibration within group in the context of machine learning. In binary classification, a score function $R(X)$ satisfies calibration within groups if $\mathrm{Pr}(Y=1\mid R(X)=r,A=a)=r$ for all score values $r$ and group level $a$, where $Y$ is the outcome variable and $X$ is the feature variables. A slightly weaker definition (Corbett-Davies et al. 2023) only requires that $Y$ is conditionally independent of $A$ given the score $R(X)$. Our fairness definition is based on this weaker definition. In KPD programs, we view $Y$ as the selection indicator and $R$ as the sensitization level of a patient. Therefore, we define our new fairness as satisfying that the selection indicator is conditionally independent of the protected feature given the sensitization level. In other words, at each sensitization level, the matching outcome is independent of the protected feature. We can view this fairness criterion in KPD programs as a reverse problem of that in machine learning. In machine learning, the randomness is due to the underlying population model, and our goal is to construct a score $R(X)$ satisfying the fairness condition. In KPD programs, we need to determine the randomness such that the observed sensitization level satisfies the fairness condition.
@@
For simplicity of presentation, we assume the protected feature $A$ is binary with two levels $\{0,1\}$ and the sensitization $R$ has $M$ levels $\{r_{1},r_{2},...,r_{M}\}$. In practice, it is common that $M=2$ where we partition all patients into highly-sensitized and lowly-sensitized patients, or $M=3$ where we partition all patients into highly-sensitized, moderately-sensitized and lowly-sensitized patients. The joint of variables $A$ and $R$ partitions all patients into $2M$ subgroups, denoted as $\{V_{ij}\}_{0\leq i\leq 1,1\leq j\leq M}$, where $V_{ij}=\{v\in V:A(v)=i,R(v)=r_{j}\}$. As in Section 2.2.2, we assign a probability distribution $\delta$ on the set of exchange plans $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Moreover, $\delta_{v}$ is the probability of each vertex $v$ to be selected in the end, and thus $V_{ij}^{-1}\sum_{v\in V_{ij}}$ is average selection rate in $V_{ij}$. Given the sensitization level $R=r_{j}$, our fairness criterion restricts the average selection rates in $V_{0j}$ and $V_{1j}$ to be close. It is natural to impose constraints that the absolute difference in these average selection rates is bound by a constant $l_{j}$. Therefore, we formulate the problem under our fairness criterion as
@@
$\max_{\delta}\ \sum_{c\in\mathcal{C}}x_{c}(\delta)u(c)\quad\text{subject to}\ \left|V_{0j}^{-1}\sum_{v\in V_{0j}}\delta_{v}-V_{1j}^{-1}\sum_{v\in V_{1j}}\delta_{v}\right|\leq l_{j},\ \forall 1\leq j\leq M.$ (4)
However, the optimization problem (4) can be difficult to solve because $\mathcal{F}$ is too large to enumerate, let alone the choice of $\delta$. We provide an approach to solve (4) based on Proposition 1 below.
Proposition 1.
There exists a solution to (4) satisfying that at most $M+1$ exchange plans in $\mathcal{F}$ have nonzero selection probability.
Instead of solving the harder optimization problem (4), by Proposition 1, we can solve an equivalent but easier mixed-integer programming:
$\max_{x_{t,c}\in\{0,1\},p_{t}\geq 0}\quad\sum_{t=1}^{M+1}\sum_{c\in\mathcal{C}}p_{t}x_{t,c}u(c)$ (5)
subject $\text{subject to}\quad&\sum_{t=1}^{M+1}p_{t}=1,&&\\ &\sum_{c\in\mathcal{C}}x_{t,c}1_{v\in c}\leq 1,\quad\forall\ v\in V,\ 1\leq t\leq M+1,\\ &|\overline{q}_{j1}-\overline{q}_{j1}|\leq l_{j},\quad\forall 1\leq j\leq M,&&\\ \end{aligned}$ (6)
where we define $\overline{q}_{ij}=|V_{ij}|^{-1}\sum_{v\in V_{ij}}\sum_{t=1}^{M+1}\sum_{c\in\mathcal{C}}p_{t}x_{t,c}1_{v\in c}$ as the average selection rate in $V_{ij}$. The optimization problem (5)–(6) involves $(M+1)|\mathcal{C}|$ binary variables and $M+1$ continuous variables. It can be efficiently solved by various optimizers, e.g., Gurobi Optimizer.
@@
For whichever choice of parameters $l_{1},...,l_{M}\geq 0$, there always exists a solution to (5)–(6). The strength of the fairness constraint depends on the parameters $l_{1},...l_{M}$. In practice, we suggest two candidate values for $l$’s: $l_{j}=1/\min\{|V_{1j}|,|V_{2j}|\}$ and $l_{j}=1/\max\{|V_{1j}|,|V_{2j}|\}$. We can interpret these two candidate values as the desired precision based on the larger or smaller subgroup. The former represents a weaker fairness constraint, and the latter represents a stronger one.

### v2

# Fairness-Aware Kidney Exchange and Kidney Paired Donation

@@ Abstract
The kidney paired donation (KPD) program provides an innovative solution to overcome incompatibility challenges in kidney transplants by matching incompatible donor-patient pairs and facilitating kidney exchanges. To address unequal access to transplant opportunities, there are two widely used fairness criteria: group fairness and individual fairness. However, these criteria do not consider protected patient features, which refer to characteristics legally or ethically recognized as needing protection from discrimination, such as race and gender. Motivated by the calibration principle in machine learning, we introduce a new fairness criterion: the matching outcome should be conditionally independent of the protected feature, given the sensitization level. We integrate this fairness criterion as a constraint within the KPD optimization framework and propose a computationally efficient solution using linearization strategies and column-generation methods. Theoretically, we analyze the associated price of fairness using random graph models. Empirically, we compare our fairness criterion with group fairness and individual fairness through both simulations and a real-data example.
Keywords: calibration, column generation, integer program, kidney paired donation, price of fairness, random graph.
@@ 1.4 Our contributions
This paper makes several contributions to the field of kidney exchange and fairness in allocation. First, we propose a new fairness criterion based on a protected feature, which has not been explored in the kidney exchange literature. We integrate this fairness criterion as a constraint within the optimization framework commonly used in kidney paired donation (KPD) programs. This flexible structure can accommodate other fairness criteria and potential recourse strategies. Furthermore, we propose a computationally efficient solution to the resulting optimization problem, achieving scalability by avoiding explicit enumeration of all exchange plans and instead relying on an iterative cutting-plane procedure.
@@ 2.1.1 Classical formulation
We represent a KPD program as a directed graph $G=(V,E)$, where $V=\{v_{1},\cdots,v_{N}\}$ denotes the vertex set and $E$ denotes the edge set. The vertex set $V$ is the set of $N$ incompatible donor-patient pairs. The edge set $E$ consists of all ordered pairs $(v_{i},v_{j})$ that the donor in vertex $v_{i}\in V$ is compatible with the patient in vertex $v_{j}\in V$. An exchange cycle, or simply cycle, is defined as a sequence of distinct incompatible donor-patient pairs. We denote a cycle $c$ as an ordered sequence of vertices $c_{1},\cdots,c_{|c|}]$ in $V$, where $|c|$ is the cycle length of $c$, satisfying that $(c_{i},c_{i+1})\in E$ for $1\leq i\leq|c|-1$, and $(c_{|c|},c_{1})\in E$. To execute the cycle, the patient in $c_{i+1}$ will receive the donor kidney of $c_{i}$ for $1\leq i\leq|c|-1$, and the patient $c_{1}$ will receive the donor kidney of $c_{|c|}$. Figure 1 illustrates how a two-way cycle and a three-way cycle work in KPD programs. An exchange plan is a collection of vertex-disjoint cycles in the graph.
@@
In fielded KPD programs, we usually restrict cycles to be no longer than 3. We use $\mathcal{C}$ to denote the set of all cycles under the restrictions on the length of the cycles. We assign a utility $u_{v_{i},v_{j}}$ to each edge $(v_{i},v_{j})\in E$. Based on the edge utility, we define the cycle utility $u(c)=\sum_{i=1}^{|c|-1}u_{c_{i},c_{i+1}}+u_{c_{|c|},c_{1}}$. In a classical KPD program, we aim to select a disjoint packing of cycles in $\mathcal{C}$ with the maximum sum of utilities, which can be formulated as the following integer program:
@@ 2.2.2 Individual fairness
The individual fairness utilizes a randomization policy to ensure that every patient has a similar chance of being matched. Let $\mathcal{F}\subseteq 2^{\mathcal{C}}$ be the set of possible exchange plans that consist of disjoint cycles in $\mathcal{C}$. We assign a probability distribution $\delta$ on $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Based on $\delta$, let $x_{c}$ be the overall probability that the cycle $c\in\mathcal{C}$ is selected as one of the disjoint subsets. Further, we can compute $\pi_{v}=\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}$, which is the probability of each vertex $v$ to be selected. Let $\pi=(\pi_{v})_{v\in V}\in\mathbb{R}^{N}$ be the vector of selection probability of all patients and let $\overline{\pi}=1^{{\mathrm{\scriptscriptstyle T}}}\pi/N$ be the average selection probability of all patients. The quantity $\pi_{v}$ is central in defining individual fairness and our new fairness criterion in Section 3.
Recall that $x_{c}$ in (1) and (2) is a binary variable indicating whether $c$ is selected or not, but here $x_{c}$ is a continuous variable bounded between 0 and 1. Following Farnadi et al. (2021), we formulate the individual fairness as a constraint in the optimization problem below
$\max_{\delta}\quad\sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\quad\sum_{v\in V}|\pi_{v}-\overline{\pi}|^{p}\geq\beta^{p}.$ (3)
Specifically, the individual fairness promotes a similar chance of being selected for each patient. We choose the $L_{p}$ norm of the vector $(\pi-\overline{\pi})$ to measure the variation of the selection probability among all the patients. We consider a fairness constraint that the $L_{p}$ norm of the vector $(\pi_{v}-\overline{\pi})$ is no less than some threshold $\beta$.
@@
3.1 Fairness formulation
In this section, we introduce a new fairness criterion based on a protected feature $A$ of patients in $V$, motivated by the concept of calibration within group in the context of machine learning. In essence, our new fairness criterion requires that the probability of a patient being selected for a transplant (the matching outcome) should not depend on the protected feature, such as race or gender, once we account for the patient’s sensitization level. This means that, within each sensitization category, patients with different protected attributes should have similar chances of being matched.
@@ 3.1 Fairness formulation
In binary classification, a score function $R(X)$ satisfies calibration within groups if $\mathrm{Pr}(Y=1\mid R(X)=r,A=a)=r$ for all score values $r$ and group level $a$, where $Y$ is the outcome variable and $X$ is the feature variables. A slightly weaker definition (Corbett-Davies et al. 2023) only requires that $Y$ is conditionally independent of $A$ given the score $R(X)$. Our fairness definition is based on this weaker definition. In KPD programs, we view $Y$ as the selection indicator and $R$ as the sensitization level of a patient. Therefore, we define our new fairness criterion as satisfying that the selection indicator is conditionally independent of the protected feature given the sensitization level. In other words, at each sensitization level, the matching outcome is independent of the protected feature. We can view this fairness criterion in KPD programs as a reverse problem of that in machine learning. In machine learning, the randomness is due to the underlying population model, and our goal is to construct a score $R(X)$ satisfying the fairness condition. In KPD programs, we need to determine the randomness such that the observed sensitization level satisfies the fairness condition.
@@
For simplicity of presentation, we assume the protected feature $A$ is binary with two levels $\{0,1\}$ and the sensitization $R$ has $M$ levels $\{r_{1},r_{2},\cdots,r_{M}\}$. In practice, it is common that $M=2$ where we partition all patients into highly-sensitized and lowly-sensitized patients, or $M=3$ where we partition all patients into highly-sensitized, moderately-sensitized and lowly-sensitized patients. The variables $A$ and $R$ partition all patients into $2M$ subgroups, denoted as $\{V_{ij}\}_{0\leq i\leq 1,1\leq j\leq M}$, where $V_{ij}=\{v\in V:A(v)=i,R(v)=r_{j}\}$. As in Section 2.2.2, we assign a probability distribution $\delta$ on the set of exchange plans $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Moreover, $\pi_{v}$ is the probability of each vertex $v$ to be selected in the end, and thus $|V_{ij}|^{-1}\sum_{v\in V_{ij}}\pi_{v}$ is average selection rate in $V_{ij}$. Given the sensitization level $R=r_{j}$, our fairness criterion restricts the average selection rates in $V_{0j}$ and $V_{1j}$ to be close. It is natural to impose constraints that the absolute difference in these average selection rates is bound by a constant $l_{j}$. Therefore, we formulate the problem under our fairness criterion as
@@
$\max_{\delta}\ \sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\ \left||V_{0j}|^{-1}\sum_{v\in V_{0j}}\pi_{v}-|V_{1j}|^{-1}\sum_{v\in V_{1j}}\pi_{v}\right|\leq l_{j},\ \forall 1\leq j\leq M.$ (4)
@@
For whichever choice of parameters $l_{1},\cdots,l_{M}\geq 0$, there always exists a solution to (4). The strength of the fairness constraint depends on the parameters $l_{1},\cdots l_{M}$. In practice, we suggest two candidate values for $l$’s: $l_{j}=0.5/\min\{|V_{1j}|,|V_{2j}|\}$ and $l_{j}=0.5/\max\{|V_{1j}|,|V_{2j}|\}$. We can interpret these two candidate values as the desired precision based on the larger or smaller subgroup. The former represents a weaker fairness constraint, and the latter represents a stronger one.
3.2 Algorithm for solving (4)
This subsection presents a linear optimization procedure for solving (4) without enumerating all exchange plans. The approach follows a master-subproblem structure similar to column generation. The master problem optimizes over the marginal probabilities of selecting each cycle and each vertex, while an auxiliary subproblem verifies whether the proposed solution is implementable and, if necessary, adds new linear constraints that refine the master problem.
Specifically, we express (4) as an optimization problem in terms of the cycle-selection variables $x_{c}$, rather than the probability distribution $\delta$ over exchange plans. Recall that the selection probability of each vertex $v$ is
$\pi_{v}=\sum_{c\in\mathcal{C}}1_{v\in c}\,x_{c}.$
Both the objective and the fairness constraints are linear in $x_{c}$. A direct constraint of variables $x_{c}$ is $0\leq x_{c}\leq 1$ for all $c\in\mathcal{C}$ and $0\leq\pi_{v}\leq 1$ for all $v\in V$. However, not every such vector $x=(x_{c})_{c\in\mathcal{C}}$ satisfying the above constraint can be realized by a valid probability distribution over disjoint exchange plans.
To enforce implementability, we introduce a scalar variable $\eta\in\mathbb{R}$ that represents the best achievable expected utility at the current selection profile and iteratively add linear inequalities, called cuts, to refine the feasible region. Let $\mathcal{B}^{\mathrm{opt}}$ and $\mathcal{B}^{\mathrm{feas}}$ denote the sets of indices for optimality and feasibility cuts generated so far; both sets start empty and expand monotonically. The restricted master problem is formulated as
$\max_{x,\pi,\eta}$ $\eta$ (MP)
subject to $\pi_{v}=\sum_{c\in\mathcal{C}}1_{v\in c}x_{c}\leq 1,\qquad\forall v\in V,$
$0\leq x_{c}\leq 1,\qquad\forall c\in\mathcal{C},$
$\left|\,|V_{0j}|^{-1}\sum_{v\in V_{0j}}\pi_{v}-|V_{1j}|^{-1}\sum_{v\in V_{1j}}\pi_{v}\,\right|\leq l_{j},\qquad j=1,\ldots,M,$
$\pi^{{\mathrm{\scriptscriptstyle T}}}\theta^{(\ell)}+\beta^{(\ell)}\geq\eta,\qquad\forall\,\ell\in\mathcal{B}^{\mathrm{opt}},$
$\pi^{{\mathrm{\scriptscriptstyle T}}}\hat{\theta}^{(k)}+\hat{\beta}^{(k)}\leq 0,\qquad\forall\,k\in\mathcal{B}^{\mathrm{feas}},$
where $(\theta^{(\ell)},\beta^{(\ell)})\in\mathbb{R}^{N}\times\mathbb{R}$ and $(\hat{\theta}^{(k)},\hat{\beta}^{(k)})\in\mathbb{R}^{N}\times\mathbb{R}$ are coefficients provided by the auxiliary subproblem.
Given a candidate $\pi$ from (MP), the auxiliary subproblem checks whether $\pi$ can be implemented by a valid distribution over exchange plans and determines the best achievable expected utility under that profile. For each plan $F_{t}\in\mathcal{F}$, let $u(F_{t})$ denote its total utility and let $\delta(F_{t})$ denote the probability that plan $F_{t}$ is selected. The auxiliary configuration problem is
$\max_{\delta}$ $\sum_{F_{t}\in\mathcal{F}}u(F_{t})\,\delta(F_{t})$ (SP)
subject to $\sum_{F_{t}\in\mathcal{F}}1_{v\in F_{t}}\,\delta(F_{t})=\pi_{v},\qquad\forall v\in V,$
$\sum_{F_{t}\in\mathcal{F}}\delta(F_{t})=1,\qquad\delta(F_{t})\geq 0,\quad\forall F_{t}\in\mathcal{F},$
with the corresponding dual formulation
$\min_{\theta\in\mathbb{R}^{N},\beta\in\mathbb{R}}$ $\pi^{{\mathrm{\scriptscriptstyle T}}}\theta+\beta$ (SP–D)
subject to $\beta+\sum_{v\in V}\theta_{v}1_{v\in F_{t}}\geq u(F_{t}),\qquad\forall F_{t}\in\mathcal{F}.$
In practice, we solve the dual problem (SP–D) rather than the primal (SP), since the primal has one variable for each exchange plan and cannot be handled directly. The dual problem is of manageable size and can be solved efficiently using column generation. The feasibility of (SP) is inferred from the behavior of (SP–D): if the dual is bounded, then (SP) is feasible, and if the dual is unbounded, then (SP) is infeasible.
[…]
The dual problem (SP–D) contains one constraint for each feasible plan $F_{t}$, which makes it impossible to construct explicitly. Column generation avoids this by starting with a small subset of constraints and iteratively adding only those that are necessary for optimality. At each iteration, the algorithm identifies the most violated dual constraint under the current dual solution $(\theta,\beta)$.
Formally, given $(\theta,\beta)$, we search for a plan $F_{t}$ whose dual constraint
$\beta+\sum_{v\in V}\theta_{v}\,1_{v\in F_{t}}\geq u(F_{t})$
is most violated. This search is equivalent to solving a maximum-weight packing problem over vertex-disjoint cycles:
$\max_{y_{c}\in\{0,1\}}\ \sum_{c\in\mathcal{C}}\Bigl(u(c)-\sum_{v\in c}\theta_{v}\Bigr)y_{c}-\beta\quad\text{subject to}\quad\sum_{c\in\mathcal{C}}y_{c}\,1_{v\in c}\leq 1,\quad\forall v\in V.$
Each binary variable $y_{c}$ indicates whether cycle $c$ is selected, and the constraint ensures that selected cycles are vertex-disjoint. If the optimal value of this problem is positive, the corresponding plan violates a dual constraint and is added to a restricted version of (SP–D). If the optimal value is nonpositive, all constraints in the full dual problem are satisfied. Because $(\theta,\beta)$ is already optimal for the restricted dual problem, and adding satisfied constraints cannot further decrease its objective value, $(\theta,\beta)$ must also be optimal for the full dual (SP–D).
The algorithm alternates between solving the master problem (MP) and invoking the auxiliary subproblem. Each iteration produces a candidate $(x,\pi,\eta)$ that satisfies the fairness constraints in (4). The auxiliary subproblem then determines whether this profile $\pi$ is implementable. If it is, the resulting dual solution provides an optimality cut that tightens the objective bound $\eta$ in (MP); otherwise, a feasibility cut removes the unattainable profile $\pi$ from the feasible region.
The process continues until no violated constraint can be found by the auxiliary subproblem. In practice, this occurs when the optimal value of the separation problem in the column-generation step is nonpositive, indicating that all dual constraints are satisfied. At this point, the restricted master problem is equivalent to the full master problem, and no further cuts or columns can improve the solution.
From a theoretical standpoint, each iteration introduces a valid linear inequality that progressively refines the feasible region, and only finitely many distinct cuts can be generated since the set of feasible plans $\mathcal{F}$ is finite. Hence, the procedure terminates after a finite number of iterations. Upon termination, the current $\pi$ is guaranteed to be implementable by a valid distribution over exchange plans, the fairness constraints are satisfied, and the corresponding objective value $\eta$ equals the optimum of (4). The final primal solution of (SP) explicitly provides this distribution, yielding an optimal randomized policy that achieves both fairness and efficiency.